# Measuring Gravitational Wave Spectrum from Electroweak Phase Transition and Higgs Self-Couplings

### v1

@@ 4.3 Likelihood
In practical analysis, the amount of data may be too large to handle directly, which necessitates the use of approximation methods. Therefore, in the literature, two approximation methods are usually considered based on the central limit theorem (CLT) and coarse graining (CG) Caprini:2024hue . In this work, we use directly the full likelihood above.
@@ 5 Fisher Information Matrix and Bayesian Analysis
In contrast, Bayesian inference, typically implemented through the MCMC sampling gelman2013bayesian , provides a more general and robust characterization of the posterior distribution. Unlike the FIM, MCMC does not assume Gaussianity or linearity, making it especially suitable for exploring non-Gaussian, multimodal, or highly degenerate likelihood surfaces. By generating a representative set of samples from the posterior distribution, MCMC yields statistically accurate estimates of parameter uncertainties and correlations. In this study, we employ both analytical derivations and simplified toy models to compare these two methods, highlighting the regimes in which the FIM offers reliable approximations and the conditions under which full Bayesian inference becomes essential.
@@ 5.1 Fisher Information Matrix and the Cramér-Rao Lower Bound
The FIM serves as a fundamental method for quantifying the amount of information that an observed random variable contains about an unknown parameter. This concept plays a crucial role in evaluating the efficiency of estimators. In particular, efficient estimators—those that achieve the minimum possible variance—are generally preferred in statistical inference. More relevantly here, the FIM is closely connected to the CRLB, which provides a theoretical lower limit on the variance of any unbiased estimator. This bound establishes a benchmark for the best achievable precision in parameter estimation, thereby offering a rigorous standard against which the performance of practical estimators can be assessed.
@@
The resulting confidence ellipse visually illustrates the uncertainties in the two parameters and their mutual correlation. A highly elongated ellipse indicates a strong correlation between the parameters, whereas a circular shape implies weak or no correlation. The shape and orientation of the ellipse thus offer valuable insights into the coupling and degeneracy structure of the parameter space, as well as the overall precision of the parameter estimation.
@@ 5.2 Bayesian Inference and MCMC Sampling
In most practical applications, particularly when dealing with high-dimensional or nonlinear models, the posterior distribution cannot be determined easily. To address this, numerical techniques are employed, among which MCMC methods are the most widely used. In this work, we perform MCMC sampling using the PyMC probabilistic programming framework AbrilPla:2023pymc , which provides efficient implementations of state-of-the-art samplers such as Metropolis and NUTS. After an initial burn-in phase, the chains are expected to adequately explore the parameter space in a manner proportional to the posterior probability density.
These methods enable the estimation of summary statistics of the posterior, such as the mean, median, credible intervals, and correlation structures between parameters. Despite its robustness, MCMC sampling can be computationally intensive, especially for complex likelihood functions or large datasets. Convergence diagnostics and autocorrelation analyses are crucial for ensuring the validity of the results. Nevertheless, when properly applied, MCMC provides an indispensable tool for Bayesian inference in both theoretical modeling and data-driven investigations.
@@ 5.3 A Toy Model : $y=kx+b$
In this subsection, we compare parameter uncertainties derived from the above two approaches. The first is based on the CRLB, which provides a theoretical minimum variance for any unbiased estimator and is computed using the FIM. The second approach employs MCMC sampling to characterize the full posterior distribution without assuming Gaussianity. Before turning to the discussion of phase transitions, we illustrate the main features of the CRLB and MCMC with a very simple example. We study a linear model that admits a straightforward analytical treatment, enabling a direct comparison of the two approaches.
@@ 6 Results
With the statistical analysis framework established in the preceding sections, we are now ready to investigate, via simulation, how the parameters in our model can be measured. This study is conducted at multiple levels of complexity. We begin with a simulation that includes only detector noises and a SGWB of astrophysical origin. Subsequently, we inject a SGWB from the sound waves, as introduced in Sec. 3. The goal is to infer the model parameters and quantify the associated uncertainties. The inferred results are then mapped onto the xSM parameter space and used to extract physical observables, such as the Higgs self-couplings.
@@ 6.1 Measurements of the Astrophysical Parameters
In contrast, the MCMC method offers a more robust and flexible framework that explores the full posterior landscape without assuming any specific distributional form. It can accurately characterize non-Gaussian features, parameter degeneracies, and multimodal distributions, albeit at the cost of higher computational demand.
@@ 6.2 Measurements of the Thermodynamics Parameters
Here, in addition to the contributions from instrumental noises and the astrophysical SGWB, we incorporate the SGWB generated by the EWPT, specifically the dominant contribution from sound waves discussed earlier. The likelihood function remains the same as in previous analyses, with a modified signal model to include this additional component:
@@
Using this extended signal model, we estimate the joint posterior distribution for six parameters: two instrumental noise parameters, $N_{\text{acc}}$ and $\delta x$; two astrophysical background parameters, $\Omega_{\text{ast}}$ and $\varepsilon$; and two parameters characterizing the phase transition SGWB, $\Omega_{0}$ and $f_{\mathrm{p}}$. The resulting confidence ellipses are shown in Fig. 6, providing a comprehensive view of the parameter correlations and uncertainties as derived from both the FIM and full posterior exploration via MCMC sampling within the Bayesian framework. This comparison allows us to assess the validity of the Gaussian approximation in a higher-dimensional, physically motivated parameter space. The injected values, used consistently in both simulation and inference, are $N_{\text{acc}}=3\times 10^{-15}$, $\delta x=8\times 10^{-12}$, $\Omega_{\text{ast}}=1\times 10^{-8}$, $\varepsilon=2/3$, $\log_{10}\Omega_{0}=-11.5$, and $\log_{10}f_{\mathrm{p}}=-2.25$. With this set of parameters, the calculated SNRs for the astrophysical background and the sound-wave contribution are 683 and 52.5, respectively. The relative uncertainties predicted by the FIM for these parameters are approximately 0.098%, 0.0069%, 0.42%, 1.38%, 0.43%, and 1.16%, respectively. In contrast, the corresponding uncertainties estimated from MCMC sampling are very similar: 0.098%, 0.0068%, 0.40%, 1.35%, 0.43%, and 1.07%.
@@
At first glance, the MCMC-derived relative uncertainties for $\Omega_{\mathrm{ast}}$ and $\varepsilon$ appear slightly smaller than those predicted by the FIM, which may seem counterintuitive since the FIM is generally expected to provide the best forecasts under the assumption of a locally Gaussian posterior. This apparent tension can be understood by recalling that Fisher forecasts probe only the local curvature of the likelihood around the maximum-likelihood point. By construction, they neglect non-Gaussian features that are naturally captured in MCMC analyses. In practice, the confidence ellipses from the FIM primarily illustrate the local shape of the posterior, whereas the standard deviations derived from MCMC chains account for the full posterior volume, including possible asymmetries, long tails, and mild non-linearities. As a result, the projected ellipses from MCMC can appear comparable or even slightly tighter, while the corresponding marginalized uncertainties remain broader. The inclusion of $\Omega_{0}$ and $f_{\mathrm{p}}$ adds further structure to the likelihood surface, since these parameters directly determine the amplitude and peak frequency of the phase-transition signal. Their coupling with $\Omega_{\text{ast}}$ and $\varepsilon$ introduces moderate degeneracies, which broaden the confidence regions along certain directions in parameter space. Overall, while the FIM remains a valuable first-order approximation, its local Gaussian nature becomes less reliable as dimensionality and model complexity increase, underscoring the need for full Bayesian inference to obtain robust uncertainty quantification.
To further investigate how the detectability of the phase transition signal depends on the fiducial model parameters, we analyze the relative uncertainty of $\log_{10}\Omega_{0}$ as a function of its injected value. In particular, we also explore how this uncertainty is influenced by the presence of an overlapping astrophysical SGWB component, characterized by different values of $\log_{10}\Omega_{\text{ast}}$. Using the FIM formalism, we compute the relative error $\frac{\Delta\log_{10}\Omega_{0}}{\log_{10}\Omega_{0}}$ under three representative scenarios: $\log_{10}\Omega_{\text{ast}}=-7.0$, 8.0-8.0, and 9.0-9.0. The results, illustrated in Fig. 7, demonstrate that a stronger astrophysical background increases the uncertainty in $\Omega_{0}$, especially when $\Omega_{0}$ is small, as expected. This highlights the importance of disentangling different SGWB components when interpreting observational data. We have also shown in this figure the comparison between MCMC-derived uncertainties (with stars) and that from the FIM for one case, which demonstrates the overall agreement between the two approaches.
@@ 6.3 Measurements of xSM Model Parameters
We then map the constrained regions of the spectral parameters to the parameter space of the xSM, characterized by the five parameters $(v_{s},m_{h_{2}},\theta,b_{3},b_{4})$, with these regions being further restricted through an extensive scan that ensures compliance with all phenomenological and theoretical constraints. For each point in this viable parameter space, we calculate the phase–transition parameters $(\alpha,\beta/H_{n},T_{n})$, and correspondingly $(\Omega_{0},f_{\mathrm{p}})$. If the prediction lies inside the FIM/MCMC acceptance bands, we keep that point. Projecting the accepted points back onto $(v_{s},m_{h_{2}},\theta,b_{3},b_{4})$ shows which parameter values are allowed and how they are related, as shown in Fig. 10. As the constraints from the FIM and the MCMC method are similar, we will now, and in the following, show only the more precise results obtained from the MCMC.
From these plots, we can see that the constrained region on the spectral parameters now leads to a similarly constrained parameter space for the particle physics model. We also note the apparent impact from the problem of parameter degeneracy which results in a more spread-out feature of the $68\%$ and $95\%$ regions. This is obvious, as the mapping
@@
is not unique. Reducing from $5$ parameters to $2$ always creates degeneracies. This means that the FIM and MCMC results do not select a single point, but instead give bands or islands of possible parameter values that can produce the same GW spectrum. The projections in Fig. 10 show these degeneracies clearly. These GW-selected regions provide a compact, data-driven summary of the xSM parameter space that is consistent with a potential detection.
@@ 6.4 Measurements of Higgs Self-Couplings
With the xSM parameter range obtained from the GW measurement of the simulated signal, we can go one step further and make predictions for various observables that may, or may not, be measured at colliders and other experiments. One particularly important set of observables of this kind is the Higgs self-couplings, or equivalently deviations of the Higgs self-couplings, $\delta\kappa_{3}$ and $\delta\kappa_{4}$, from their SM values, as defined in Eq. (2.8). These two couplings describe the shape of the scalar potential, and are therefore important for understanding the electroweak symmetry breaking and for exploring possible extensions of the Higgs sector.
@@
The question naturally arises as to whether the GW measurement from EWPT can be used to give a better determination of these two couplings. To do this, from the parameter space of the xSM identified from the GW measurement, we infer the corresponding $\delta\kappa_{3}$ and $\delta\kappa_{4}$. These values are added onto Fig. 11, where the blue points denote all the points that can give a first-order EWPT and survive phenomenological constraints. The green and red points denote those that give the GW measurement within one and two standard deviations, respectively, the same as in previous sections. In this way, we are actually combining future GW measurements with current collider and other phenomenological and theoretical constraints.
@@
For all points, due to features of this model, as explained in more detail in Alves:2018oct , there appears a linear correlation between these two couplings, which is more easily understood from a Taylor expansion in powers of the mixing angle $\theta$, taking into account also the concentrated regions of the parameter space resulting from all phenomenological and theoretical constraints. Comparing the blue with the green and red points, we can see that the region on the plane $(\delta\kappa_{3},\delta\kappa_{4})$ is narrowed down to a much smaller one due to the GW measurement. This highlights the importance of GW measurement in probing the Higgs couplings, especially regarding the quartic ones. We note, however, that due to the same parameter degeneracy problem, the precision on $\delta\kappa_{3}$ and $\delta\kappa_{4}$ is limited, a problem at least for this model, which could potentially affect a much broader class of models. A potential solution to this conundrum is to make use of finer structures on the spectrum, such as the damping feature at the higher frequency caused by dissipative effects in the fluid Guo:2023koq . We further note that in deriving the results here, we have neglected theoretical uncertainties coming from spectrum and phase transition parameter calculations, as explored recently Lewicki:2024xan based on the Fisher information method, the inclusion of which would lead to a less constraining result. Systematically quantifying and including these uncertainties in the Bayesian framework and with a more faithful detector simulation would be a direction for studies in the future.
@@ 7 Conclusion
In this work, we have developed a comprehensive framework that connects theoretical modeling of FOPTs in the xSM with realistic data analysis strategies for space-based GW detectors such as Taiji and LISA. By combining frequency-domain detector response modeling, astrophysical foregrounds, and instrumental noise, we constructed simulated datasets for the SGWB and carried out both FIM and Bayesian MCMC analyses. Using simulated data, we first estimated instrumental and astrophysical parameters and then extended the analysis to include the stochastic signal from sound waves generated during a first-order EWPT. The inferred spectral parameters $(\Omega_{0},f_{\mathrm{p}})$ were subsequently mapped onto the xSM parameter space $(v_{s},m_{h_{2}},\theta,b_{3},b_{4})$, enabling us to identify viable regions consistent with a potential SGWB detection. This mapping further allowed us to translate GW constraints into predictions for Higgs self-coupling deviations $(\delta\kappa_{3},\delta\kappa_{4})$, which encode the shape of the electroweak scalar potential. Our analysis shows that space-based GW observations can help determine these couplings. While collider measurements of $\delta\kappa_{3}$ and especially $\delta\kappa_{4}$ remain extremely challenging, GW-based inference from FOPTs offers an indirect but powerful probe of scalar self-interactions. The synergy between collider physics and GW astronomy thus opens a new pathway toward a more complete understanding of electroweak symmetry breaking and the origin of the Higgs potential. Future work will extend this framework to include detector networks, realistic data gaps, and a joint analysis within the global-fit approach that simultaneously accounts for deterministic sources such as massive black-hole binaries. The inclusion of theoretical uncertainties in the phase-transition modeling—such as gauge dependence and finite-temperature corrections—will also be an important next step toward robust, data-driven constraints on beyond-Standard-Model physics from the upcoming era of space-based GW observations.

### v2

@@ 4.3 Likelihood
In practical analysis, the amount of data may be too large to handle directly, which necessitates the use of approximation methods. Therefore, in the literature, two approximation methods are usually considered based on the central limit theorem (CLT) and coarse graining (CG) Caprini:2024hue. In this work, we use the full likelihood above directly.
@@ 5 Fisher Information Matrix and Bayesian Analysis
In contrast, Bayesian inference, typically implemented through MCMC sampling gelman2013bayesian, provides a more general and robust characterization of the posterior distribution. Unlike the FIM, MCMC does not assume Gaussianity or linearity, making it especially suitable for exploring non-Gaussian, multimodal, or highly degenerate likelihood surfaces. By generating a representative set of samples from the posterior distribution, MCMC yields statistically accurate estimates of parameter uncertainties and correlations. In this study, we employ both analytical derivations and simplified toy models to compare these two methods, highlighting the regimes in which the FIM offers reliable approximations and the conditions under which full Bayesian inference becomes essential.
@@ 5.1 Fisher Information Matrix and the Cramér-Rao Lower Bound
The FIM serves as a fundamental method for quantifying the amount of information that an observed random variable contains about an unknown parameter. This concept plays a crucial role in evaluating the efficiency of estimators. In particular, efficient estimators—those that achieve the minimum possible variance—are generally preferred in statistical inference. More relevantly, the FIM is closely connected to the CRLB, which provides a theoretical lower limit on the variance of any unbiased estimator. This bound establishes a benchmark for the best achievable precision in parameter estimation, thereby offering a rigorous standard against which the performance of practical estimators can be assessed.
@@
The resulting confidence ellipse visually illustrates the uncertainties in the two parameters and their mutual correlation. A highly elongated ellipse indicates a strong correlation between the parameters, whereas a circular shape implies a weak or no correlation. The shape and orientation of the ellipse thus offer valuable insights into the coupling and degeneracy structure of the parameter space, as well as the overall precision of the parameter estimation.
@@ 5.2 Bayesian Inference and MCMC Sampling
In most practical applications, particularly when dealing with high-dimensional or nonlinear models, the posterior distribution cannot be easily determined. To address this, numerical techniques are employed, among which MCMC methods are the most widely used. In this work, we perform MCMC sampling using the PyMC probabilistic programming framework AbrilPla:2023pymc, which provides efficient implementations of state-of-the-art samplers such as Metropolis and NUTS. After an initial burn-in phase, the chains are expected to explore the parameter space adequately in a manner proportional to the posterior probability density.
These methods enable the estimation of summary statistics of the posterior, such as the mean, median, credible intervals, and correlation structures between parameters. Despite its robustness, MCMC sampling can be computationally intensive, especially for complex likelihood functions or large datasets. Convergence diagnostics and autocorrelation analyzes are crucial for ensuring the validity of the results. Nevertheless, when properly applied, MCMC provides an indispensable tool for Bayesian inference in both theoretical modeling and data-driven investigations.
@@ 5.3 A Toy Model : $y=kx+b$
In this subsection, we compare the parameter uncertainties derived from the two approaches mentioned above. The first is based on the CRLB, which provides a theoretical minimum variance for any unbiased estimator and is computed using the FIM. The second approach employs MCMC sampling to characterize the full posterior distribution without assuming Gaussianity. Before turning to the discussion of phase transitions, we illustrate the main features of the CRLB and MCMC with a very simple example. We study a linear model that admits a straightforward analytical treatment, enabling a direct comparison of the two approaches.
@@ 6 Results
With the statistical analysis framework established in the preceding sections, we are now ready to investigate, via simulation, how the parameters in our model can be measured. This study is conducted at multiple levels of complexity. We begin with a simulation that includes only detector noise and a SGWB of astrophysical origin. Subsequently, we inject a SGWB from the sound waves, as introduced in Sec. 3. The goal is to infer the model parameters and quantify the associated uncertainties. The inferred results are then mapped onto the xSM parameter space and used to extract physical observables, such as the Higgs self-couplings.
@@ 6.1 Measurements of the Astrophysical Parameters
In contrast, the MCMC method offers a more robust and flexible framework that explores the full posterior landscape without assuming any specific distributional form. It can accurately characterize non-Gaussian features, parameter degeneracies, and multimodal distributions, albeit at the cost of higher computational demands.
@@ 6.2 Measurements of the Thermodynamics Parameters
Here, in addition to the contributions from instrumental noises and the astrophysical SGWB, we incorporate the SGWB generated by the EWPT, specifically the dominant contribution from sound waves discussed earlier. The likelihood function remains the same as in previous analyzes, with a modified signal model to include this additional component:
@@
Using this extended signal model, we estimate the joint posterior distribution for six parameters: two instrumental noise parameters, $N_{\text{acc}}$ and $\delta x$; two astrophysical background parameters, $\Omega_{\text{ast}}$ and $\varepsilon$; and two parameters characterizing the phase transition of the SGWB, $\Omega_{0}$ and $f_{\mathrm{p}}$. The resulting confidence ellipses are shown in Fig. 6, providing a comprehensive view of the parameter correlations and uncertainties as derived from both the FIM and full posterior exploration via MCMC sampling within the Bayesian framework. This comparison allows us to assess the validity of the Gaussian approximation in a higher-dimensional, physically motivated parameter space. The injected values, used consistently in both simulation and inference, are $N_{\text{acc}}=3\times 10^{-15}$, $\delta x=8\times 10^{-12}$, $\Omega_{\text{ast}}=1\times 10^{-8}$, $\varepsilon=2/3$, $\log_{10}\Omega_{0}=-11.5$, and $\log_{10}f_{\mathrm{p}}=-2.25$. With this set of parameters, the calculated SNRs for the astrophysical background and the sound-wave contribution are 683 and 52.5, respectively. The relative uncertainties predicted by the FIM for these parameters are approximately 0.098%, 0.0069%, 0.42%, 1.38%, 0.43%, and 1.16%, respectively. In contrast, the corresponding uncertainties estimated from MCMC sampling are very similar: 0.098%, 0.0068%, 0.40%, 1.35%, 0.43%, and 1.07%.
@@
At first glance, the MCMC-derived relative uncertainties for $\Omega_{\mathrm{ast}}$ and $\varepsilon$ appear slightly smaller than those predicted by the FIM, which may seem counterintuitive since the FIM is generally expected to provide the best forecasts under the assumption of a locally Gaussian posterior. This apparent tension can be understood by recalling that Fisher forecasts probe only the local curvature of the likelihood around the maximum-likelihood point. By construction, they neglect non-Gaussian features that are naturally captured in MCMC analyzes. In practice, the confidence ellipses from the FIM primarily illustrate the local shape of the posterior, whereas the standard deviations derived from MCMC chains account for the full posterior volume, including possible asymmetries, long tails, and mild non-linearities. As a result, the projected ellipses from MCMC can appear comparable or even slightly tighter, while the corresponding marginalized uncertainties remain broader. The inclusion of $\Omega_{0}$ and $f_{\mathrm{p}}$ adds further structure to the likelihood surface since these parameters directly determine the amplitude and peak frequency of the phase-transition signal. Their coupling with $\Omega_{\text{ast}}$ and $\varepsilon$ introduces moderate degeneracies, which broaden the confidence regions along certain directions in parameter space. Overall, while the FIM remains a valuable first-order approximation, its local Gaussian nature becomes less reliable as dimensionality and model complexity increase, underscoring the need for full Bayesian inference to obtain robust uncertainty quantification.
To further investigate how the detectability of the phase transition signal depends on the fiducial model parameters, we analyze the relative uncertainty of $\log_{10}\Omega_{0}$ as a function of its injected value. In particular, we also explore how this uncertainty is influenced by the presence of an overlapping astrophysical SGWB component, characterized by different values of $\log_{10}\Omega_{\text{ast}}$. Using the FIM formalism, we compute the relative error $\frac{\Delta\log_{10}\Omega_{0}}{\log_{10}\Omega_{0}}$ under three representative scenarios: $\log_{10}\Omega_{\text{ast}}=-7.0$, 8.0-8.0, and 9.0-9.0. The results, illustrated in Fig. 7, demonstrate that a stronger astrophysical background increases the uncertainty in $\Omega_{0}$, especially when $\Omega_{0}$ is small, as expected. This highlights the importance of disentangling different SGWB components when interpreting observational data. We have also shown in this figure the comparison between MCMC-derived uncertainties (with stars) and those from the FIM for one case, which demonstrates the overall agreement between the two approaches.
@@ 6.3 Measurements of xSM Model Parameters
We then map the constrained regions of the spectral parameters to the parameter space of the xSM, characterized by the five parameters $(v_{s},m_{h_{2}},\theta,b_{3},b_{4})$. These regions are further restricted through an extensive scan that ensures compliance with all phenomenological and theoretical constraints. For each point in this viable parameter space, we calculate the phase–transition parameters $(\alpha,\beta/H_{n},T_{n})$ and correspondingly $(\Omega_{0},f_{\mathrm{p}})$. If the prediction lies within the FIM/MCMC acceptance bands, we will keep that point. Projecting the accepted points back onto $(v_{s},m_{h_{2}},\theta,b_{3},b_{4})$ shows which parameter values are allowed and how they are related, as shown in Fig. 10. As the constraints from the FIM and the MCMC method are similar, we will now, and in the following discussion, show only the more precise results obtained from the MCMC.
From these plots, we can see that the constrained region on the spectral parameters now leads to a similarly constrained parameter space for the particle physics model. We also note the apparent impact of the problem of parameter degeneracy, which results in a more spread-out feature of the $68\%$ and $95\%$ regions. This is obvious, as the mapping
@@
is not unique. Reducing from $5$ parameters to $2$ always creates degeneracies. This means that the FIM and MCMC results do not select a single point; instead, they provide bands or islands of possible parameter values that can produce the same GW spectrum. The projections in Fig. 10 show these degeneracies clearly. These GW-selected regions provide a compact, data-driven summary of the xSM parameter space that is consistent with a potential detection.
@@ 6.4 Measurements of Higgs Self-Couplings
With the xSM parameter range obtained from the GW measurement of the simulated signal, we can take one step further and make predictions for various observables that may or may not be measured at colliders and other experiments. One particularly important set of observables of this kind is the Higgs self-couplings, or equivalently, deviations of the Higgs self-couplings, $\delta\kappa_{3}$ and $\delta\kappa_{4}$, from their SM values, as defined in Eq. (2.8). These two couplings describe the shape of the scalar potential and are therefore important for understanding electroweak symmetry breaking and for exploring possible extensions of the Higgs sector.
@@
The question naturally arises as to whether the GW measurement from EWPT can be used to provide a better determination of these two couplings. To do this, from the parameter space of the xSM identified from the GW measurement, we infer the corresponding $\delta\kappa_{3}$ and $\delta\kappa_{4}$. These values are added onto Fig. 11, where the blue points denote all the points that can give a first-order EWPT and survive phenomenological constraints. The green and red points denote those that provide the GW measurement within one and two standard deviations, respectively, as in previous sections. In this way, we are actually combining future GW measurements with current collider and other phenomenological and theoretical constraints.
@@
For all points, due to the features of this model, as explained in more detail in Alves:2018oct, there appears to be a linear correlation between these two couplings, which can be more easily understood from a Taylor expansion in powers of the mixing angle $\theta$, taking into account the concentrated regions of the parameter space resulting from all phenomenological and theoretical constraints. Comparing the blue points with the green and red points, we can see that the region on the plane $(\delta\kappa_{3},\delta\kappa_{4})$ has narrowed down to a much smaller area due to the GW measurement. This highlights the importance of GW measurement in probing the Higgs couplings, especially regarding the quartic ones. We note, however, that due to the same parameter degeneracy problem, the precision on $\delta\kappa_{3}$ and $\delta\kappa_{4}$ is limited. This limitation, at least for this model, could potentially affect a much broader class of models. A potential solution to this conundrum is to make use of finer structures on the spectrum, such as the damping feature at higher frequencies caused by dissipative effects in the fluid Guo:2023koq. We further note that in deriving the results here, we have neglected theoretical uncertainties arising from spectrum and phase transition parameter calculations, as explored recently Lewicki:2024xan. The inclusion of these would lead to a less constraining result. Systematically quantifying and including these uncertainties in the Bayesian framework, along with a more faithful detector simulation, would be a direction for future studies.
@@ 7 Conclusion
In this work, we have developed a comprehensive framework that connects the theoretical modeling of FOPTs in the xSM with realistic data analysis strategies for space-based GW detectors, such as Taiji and LISA. By combining frequency-domain detector response modeling, astrophysical foregrounds, and instrumental noise, we constructed simulated datasets for the SGWB and carried out both FIM and Bayesian MCMC analyzes. Using simulated data, we first estimated instrumental and astrophysical parameters, and then extended the analysis to include the stochastic signal from sound waves generated during a first-order EWPT. The inferred spectral parameters $(\Omega_{0},f_{\mathrm{p}})$ were subsequently mapped onto the xSM parameter space $(v_{s},m_{h_{2}},\theta,b_{3},b_{4})$, enabling us to identify viable regions consistent with a potential SGWB detection. This mapping further allowed us to translate GW constraints into predictions for Higgs self-coupling deviations $(\delta\kappa_{3},\delta\kappa_{4})$, which encode the shape of the electroweak scalar potential. Our analysis shows that space-based GW observations can help determine these couplings. While collider measurements of $\delta\kappa_{3}$ and especially $\delta\kappa_{4}$ remain extremely challenging, GW-based inference from FOPTs offers an indirect but powerful probe of scalar self-interactions. The synergy between collider physics and GW astronomy thus opens a new pathway toward a more complete understanding of electroweak symmetry breaking and the origin of the Higgs potential. Future work will extend this framework to include detector networks, realistic data gaps, and a joint analysis within the global-fit approach that simultaneously accounts for deterministic sources such as massive black-hole binaries. The inclusion of theoretical uncertainties in the phase-transition modeling—such as gauge dependence and finite-temperature corrections—will also be an important next step toward robust, data-driven constraints on beyond-Standard-Model physics from the upcoming era of space-based GW observations.